How do I solve this to get y=mx+b
x-2y=1 please help thanks

Answer by lenny460(1073)

How do I solve this to get y=mx+b
x-2y=1 please help thanks

x - 2y = 1
Subtract x from each side of the equation.
x - 2y - x = 1 - x
Combine like terms:
x - x = 0
-2y = 1 - x
Divide each item by -2
-2y/-2 = 1/-2 -x/-2
Answer:
y = -1/2 + x/2

With the above now you can find the coordinates and graph.